OpenStudy (nuccioreggie):

Solve the equation x divided by six point eight equals eleven point five. x equals what value? negative seventy-eight point two negative one point seven one point seven seventy-eight point two

OpenStudy (nuccioreggie):

@mathmath333

OpenStudy (mathmath333):

can u form the equation given in the question

OpenStudy (nuccioreggie):

i think is its c but yea

OpenStudy (anonymous):

6.8*11.5= 78.2

OpenStudy (anonymous):

Where did you get 7.8 mathmath333?

OpenStudy (mathmath333):

oh typo \(\dfrac{x}{6.8}=11.5\)

OpenStudy (anonymous):

You have your answer. X/6.8= 11.5 6.8*11.5= 78.2